OpenStudy (anonymous):

you, in your new red Porsche, decide to make a 160km trip to Carolina Beach, traveling at 80km?hr. Your make a return trip at a rate of 48km?hr. What was your average speed for the entire trip?

OpenStudy (anonymous):

figure out the total time, and the total distance divide the total distance by the total time

OpenStudy (anonymous):

ok it was easy but i got it wrong total distance is \(2\times 160=320\)

OpenStudy (anonymous):

time going in your new red porche is \(\frac{160}{80}=2\)

OpenStudy (anonymous):

how about the time returning?

OpenStudy (anonymous):

ugh i don't know cause it comes out as a decimal?

OpenStudy (anonymous):

much easier (believe it or not) if you leave it as a fraction

OpenStudy (anonymous):

in other words, take \(\frac{160}{48}\) and instead of writing it as a repeating decimal, reduce it to \(\frac{10}{3}\)

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

okay so? then do i leave the other as 2? or 160/2?

OpenStudy (anonymous):

no use 2

OpenStudy (anonymous):

now the next job is to total up the time, in other words compute \[2+\frac{10}{3}\] leave it as a fraction

OpenStudy (anonymous):

okay then what?

OpenStudy (anonymous):

16/3

OpenStudy (anonymous):

right and now speed is distance divided by rate

OpenStudy (anonymous):

so your final job is \[320\div \frac{16}{3}\]

OpenStudy (anonymous):

So the average is 60?

OpenStudy (anonymous):

that is what i get too

OpenStudy (anonymous):

kay good. Thank you!!

OpenStudy (anonymous):

notice two things 1) the average speed is not the "average of the speeds" and 2) the 160 is a red herring if you really want to impress your math teacher, show that it works no matter how far you drive

OpenStudy (anonymous):

you could drive 160 miles or 100 miles or 3,000 miles or \(m\) miles, the average will still be 60

OpenStudy (anonymous):

Okay Thank you soo much. =)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

try it with \(m\) you get \[\frac{2m}{\frac{m}{80}+\frac{m}{48}}=60\] after some algebra